OpenStudy (anonymous):

How would I go about finding the volume of 0.100 M hydrochloric acid [HCl] necessary to react completely with 1.51 grams of Aluminum Hydroxide [Al(OH)3]?

OpenStudy (greatlife44):

first set up the balanced equation

OpenStudy (anonymous):

Yup, I got that part already. I'm need the help with the actual formulas and their sequencing in hopes to find the correct answer.

OpenStudy (greatlife44):

okay, let me see

OpenStudy (greatlife44):

\[Al(OH)_{3} + 3HCl --> AlCl_{3} + 3H2O\]

OpenStudy (anonymous):

That's the same balanced equation I concluded with, too.

OpenStudy (anonymous):

Now, to use the correct formulas and then plugging in the numbers/info...

OpenStudy (greatlife44):

okay say we have 1.51 grams of aluminum hydroxide right? we convert aluminum hydroxide to moles. \[1.51 grams Al(OH)_{3}*(\frac{ mol }{ 78.6g }) = 0.019 moles Al(OH)_{3}\] then we need to find how moles of HCL we need. so we use the molar ratio. this is from our reaction \[0.019 mol Al(OH)_{3}*(\frac{ 3HCL }{ Al(OH)_{3}}) = 0.057 moles of HCL\] \[M_{a}V_{a} = M_{B}v_{B}\] But we know already the molarity of HCL and we've found how many moles react. so all can do is this. \[Molarity = \frac{ mol }{ L }\] \[\frac{ 0.1 mol }{ L } = \frac{ 0.057moles }{ x } = 0.57L \]

OpenStudy (greatlife44):

essentially, this is the power of using the balanced equation and stoichiometry

OpenStudy (anonymous):

So earlier today, when I had just begun trying to find the answer to this question, I initially began to use: \[M _{1}V_{1}=M_{2}V_{2}\] but then for some reason was having trouble solving so I assumed I was using the incorrect formula and started over so that I could go about solving it a completely different way...so I was right about the formula from the beginning and shouldn't have second guessed myself then, right?

OpenStudy (greatlife44):

I'm thinking that for this you need to go back and find how many moles of each reactant that you would need to react in order to get the reaction going to completion.

OpenStudy (greatlife44):

you don't know the molarity of Aluminum hydroxide or the volume.

OpenStudy (greatlife44):

so the best option here would be to find the number of moles of HCL you need and then figure out how many mL you need which for that you can do because in the equation you know the molarity for HCL. in that case, you have only 1 unknown. but if you use the molarity equation you have 2 unknowns. I guess that's why I used the method above.